# Shen2018S18

### v1

@@ 1 Introduction
Several parameters quantify the goodness of fit of Gaia five-parameter solutions. The most useful of these is RUWE (“renormalized unit weight error”) which approximately represents the reduced chi-squared, $\chi_{\nu}^{2}$, of the five-parameter astrometric solution (Lindegren, 2018). Several previous works have explored the utility of RUWE as diagnostic of spurious solutions and binary companions. Lindegren (2018) proposed that for Gaia DR2, a cut of RUWE $<1.4$ would retain most “good” solutions while filtering out most spurious ones. [^1] For DR3, a threshold of RUWE $>1.25$ is more appropriate (Penoyre et al., 2022a). Belokurov et al. (2020) showed that, at least for nearby sources, RUWE $>1.4$ primarily selects sources in regions of the color-magnitude diagram (CMD) where binaries are expected to be overrepresented. Penoyre et al. (2020) developed a framework to predict RUWE and other astrometric observables for binaries; Penoyre et al. (2022a), Penoyre et al. (2022b), and Korol et al. (2022) used this framework to infer properties of the local binary population from the observed RUWE distribution. Most recently, Castro-Ginard et al. (2024) developed a model to quantify the selection function of RUWE-based binary samples, while Kiefer et al. (2024) used RUWE to select candidate exoplanet hosts.
It is frequently assumed in the literature that parallaxes for source with large RUWE (most commonly, RUWE $>1.4$) are unreliable. However, it has not been quantified how unreliable, and under what conditions such parallaxes still contain useful information. One might expect, for example, that larger RUWE values result in less reliable parallaxes, or that nearby sources with large RUWE are less affected than distant ones with the same RUWE. Sources with elevated RUWE are often of particular astrophysical interest: for example, binaries that receive spectroscopic solutions but not astrometric orbital solutions are likely to have large RUWE (e.g. Thompson et al., 2019; Yamaguchi et al., 2024; Wang et al., 2024), and distance is often the key uncertainty in measuring the physical properties of these systems.
@@ 2.2 Validating RUWE predictions from gaiamock
gaiamock predicts RUWE for DR3 binaries quite accurately, with negligible systemic offset and scatter of 16.6%. RUWE is strongly correlated with $a_{0}$, which is not surprising, because it is essentially a measure of the size of typical residuals for the five-parameter solution (which scales with $a_{0}$) in units of the per-epoch astrometric uncertainties. However, predicting RUWE with gaiamock results in a tighter correlation than any simple scaling with $a_{0}$. The tight correlation shown in Figure 1 demonstrates that gaiamock predicts RUWE values for binaries robustly.
@@ 5 Discussion and Conclusions
Our RUWE-based parallax inflation factors can supersede previously published prescriptions for sources with “good” parallaxes. For example, El-Badry et al. (2021) found that the parallax uncertainties of sources with RUWE $<1.4$ must be inflated by a magnitude-dependent factor to achieve statistical consistency between the parallaxes of the components of wide binaries, and should be inflated by larger factors for sources with close neighbors. Our modeling in this work suggests that those inflation factors were driven mainly by sources with RUWE between 1 and 1.4, some of which still have imperfect astrometric fits.
@@ acknowledgments
We thank Soumyadeep Bhattacharjee for discussions that prompted this work. This research was supported by NSF grant AST-2307232 and donations from Isaac Malsky. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.

### v2

@@ 1 Introduction
Several parameters quantify the goodness of fit of Gaia five-parameter solutions. The most useful of these is RUWE (“renormalized unit weight error”) which approximately represents the reduced chi-squared, $\chi_{\nu}^{2}$, of the five-parameter astrometric solution (Lindegren, 2018). Several previous works have explored the utility of RUWE as diagnostic of spurious solutions and binary companions. Lindegren (2018) proposed that for Gaia DR2, a cut of RUWE $<1.4$ would retain most “good” solutions while filtering out most spurious ones. [^1] For DR3, a threshold of RUWE $<1.25$ has been proposed to be more appropriate, though the exact threshold varies with sky position (Penoyre et al., 2022a). Belokurov et al. (2020) showed that, at least for nearby sources, RUWE $>1.4$ primarily selects sources in regions of the color-magnitude diagram (CMD) where binaries are expected to be overrepresented. Penoyre et al. (2020) developed a framework to predict RUWE and other astrometric observables for binaries; Penoyre et al. (2022a), Penoyre et al. (2022b), and Korol et al. (2022) used this framework to infer properties of the local binary population from the observed RUWE distribution. Most recently, Castro-Ginard et al. (2024) developed a model to quantify the selection function of RUWE-based binary samples, while Kiefer et al. (2024) used RUWE to select candidate exoplanet hosts.
It is frequently assumed in the literature that parallaxes for source with large RUWE (most commonly, RUWE $>1.4$) are unreliable. However, it has not been quantified how unreliable, and under what conditions such parallaxes still contain useful information. One might expect, for example, that larger RUWE values result in less reliable parallaxes, or that nearby sources with large RUWE are less affected than distant ones with the same RUWE. Sources with elevated RUWE are often of particular astrophysical interest: for example, binaries that receive spectroscopic solutions but not astrometric orbital solutions are likely to have large RUWE (e.g. Thompson et al., 2019; El-Badry & Quataert, 2021; Yamaguchi et al., 2024; Wang et al., 2024), and distance is often the key uncertainty in measuring the physical properties of these systems.
@@ 2.2 Validating RUWE predictions from gaiamock
gaiamock predicts RUWE for DR3 binaries quite accurately, with negligible systemic offset and scatter of 16.6%. RUWE is strongly correlated with $a_{0}$, which is not surprising, because it is essentially a measure of the size of typical residuals for the five-parameter solution (which scales with $a_{0}$) in units of the per-epoch astrometric uncertainties. However, predicting RUWE with gaiamock results in a tighter correlation than any simple scaling with $a_{0}$. The tight correlation shown in Figure 1 demonstrates that gaiamock predicts RUWE values for binaries robustly. At low RUWE, there is a slight bias toward higher predicted than observed RUWE, perhaps hinting at imperfectly modeled systematics in the low-RUWE regime. This, however, only amounts to a median ruwe difference of 0.02 for sources with RUWE $<1.5$.
@@ 5 Discussion and Conclusions
While our fiducial simulations assume that elevated RUWE is always a consequence of binarity, we have verified that our inferred uncertainty inflation factors still perform well when we simply add unmodeled noise to the epoch astrometry, as might be expected if an observed source’s RUWE is a result of photometric variability or unmodeled instrumental issues.
5.1 Uncertainties for sources with low RUWE
El-Badry et al. (2021) found that the parallax uncertainties of sources with RUWE $<1.4$ must be inflated by a magnitude-dependent factor to achieve statistical consistency between the parallaxes of the components of wide binaries. We find that the uncertainty inflation factors inferred by El-Badry et al. (2021) are nearly unchanged when a cut of RUWE $<1.1$ is used instead of RUWE $<1.4$ and conclude that some uncertainty inflation is still required for sources with RUWE $\approx 1$. That is, the inflation factors inferred by El-Badry et al. (2021) are not simply a result of sources with e.g. RUWE $=1.2-1.4$. We advocate using the uncertainty inflation factors inferred here for sources with significantly inflated RUWE, and reverting to the prescriptions from El-Badry et al. (2021) for other sources.
A reasonable sky-averaged RUWE threshold for poor astrometric fits, above which uncertainties should be inflated following Equation 3, is RUWE $>1.25$ (e.g. Penoyre et al., 2022a). The distribution of RUWE for well-behaved sources varies somewhat across the sky, so another option is to identify sources with poor fits as those with RUWE significantly larger than average for their sky position and apparent magnitude. However, the origin of spatial trends in RUWE is not yet well understood, so caution is advisable in modeling them. We also not advocate deflating uncertainties for sources with RUWE $<1$: our empirical tests with wide binaries find that such sources have similarly true uncertainties to sources with RUWE $\approx 1$.
@@ acknowledgments
We thank the referee for helpful comments and Soumyadeep Bhattacharjee for discussions that prompted this work. This research was supported by NSF grant AST-2307232 and donations from Isaac Malsky. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.